OpenStudy (anonymous):

A quadratic function has its vertex at the point (2,3). The function passes through the point (3,9). Find the quadratic and linear coefficients and the constant term of the function

OpenStudy (anonymous):

I found the quadratic coefficient is 6 but my other answers were wrong

OpenStudy (anonymous):

The function is y = x^2 -5x +6

OpenStudy (anonymous):

th quadratic coefficient is 1 and the linear coefficient is -5

OpenStudy (anonymous):

A quadratic can be written as \(f(x) =a(x-h)^{2} + k\) where (h,k) is the vertex. So it tells you your vertex is (2,3) and that the point (3,9) exists. So h must be 2 and k must be 3, giving you \(f(x) = a(x-2)^{2} + 3\) Now the second point says that when x is 3, f(x) = 9. So now just plug in 3 for x and 9 for f(x) and solve for a: \(9=a(3-2)^{2} + 3\)